A Complete Bibliography of Publications in Biometrika for the decade 1900–1909

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.04

Title word cross-reference

1 [Eld03c]. 100 [Eld03c]. A [Pea09c]. B [Pea09c]. Γ [Duf09]. Γ(x + 1) [Ano08g]. n [Eld08]. z = f(x_{1}, x_{2}) [Pea09a].

&c [Lat02].

-Dimension [Eld08]. -Function [Duf09].

1895 [PY02]. 1907 [E.07g].

Abnormal [Dav02]. Actinosphaerium [Smi03]. Acute [Bro07]. Adams [P.09]. Addendum [Eld03a, Lew08a]. Adoxa [Har09b]. Adverse [E.09b]. After [Pea06b]. Age [Ano04c]. Albinism [Bat04, Bat05, W.04]. Albino [Dar02, Dar03a, Dar03b, Dar04]. Albinos [Dar03b, Sch05b]. Allied [Ano04b]. Alternative [Wel02a]. Ambiguity [Wel02b]. America [Sch06a]. among [E.09b]. Analysis [Eld03a, Eld03b, Gal09]. Anatomy [Sch05a]. Ancestral [P.04a, PYBL03, Sch06b]. Anemone [Yul02b]. Angaben [P.07b]. Anthropometric [Toc07, Pow01]. Anthropometry [Mac02a]. Anti [May09]. Anti-Typhoid [May09]. antisymmetry [Dun03]. any [Ano08c]. Aphis [War02]. Apparent [Ano02b]. Appendix [Toc06a, Toc06b]. Arbustorum [Di 07]. Arcella [PD03]. Aristata [Rob08]. Arithmetical [Pea03b]. Arrangement [Bry02, Pea06f]. Array [Gal07]. Article [E.07g]. Ascertain [Lat05, War03]. Association [Eld09, Yul03]. Assortative [Ano03a]. Aster [W.02]. Asylums [Ano06f, Toc07]. Asymmetrie [Dun03]. Atra [Edw08]. Attack [Pea03b, Pea06h, Tur07]. Attempt [Lat05, War03]. Attributes [Yul03]. Auratus [Kri08]. Aurel [P.03b, Pea03b]. Aurelia [Bro01]. Aurita [Bro01]. Australia [Pow01].

B [Ano08d]. Barrow [Sch05a]. Based [War09]. Basis [P.07a]. Bateson [Wel03]. Bees [Edg07]. Between [Ano04c, Dar03b, Jen06, P.04b, Pea09c, Pea09b, Per04b, S02, Tsc05b, Tur06, War03, Har09a, Jen09]. Bibliography [Ano06a, Ano07, Ano08a, Ano09a]. Binary [Sim02]. Biology [Pea02d]. Biometric [BLP05, GW09, Pea09d, Gai04, Hen02]. Biometrical [Pea05a, Pea07a, Smi03]. Biometrika [Ano01a, Ano01b, Ano09c]. Biométrique [Gai04]. Biometrische [Hen02]. Biometry [E.07g, Gal01]. Birth [Pea06b, Ano08d]. Blattidae [Sch08, Sch08]. Blood [Pea09d]. Body [War03]. Bony [Bry02]. Book [Ano08e, Ano08d, Ano08c, Ano08b, Ano09b, E.06, E.07e, E.07a, E.07c, E.07f, E.07b, E.07d, E.08, E.09b, E.09a, Gor09, P.07a, P.07b, P.09]. Boyle [E.08]. Brachycephalie [P.07b]. Brachydactyly [Lew08b]. Brain [BLP05, Gla05, Pea05a]. Brain-Weight [Pea05a]. Brain-Weights [BLP05]. Break [Lat05]. Breeding [Gal09, Woo03, E.09a]. Bureau [E.07g]. Butter [Rie09].

C [E.06, P.03a]. Calculating [Har09c]. Calculation [Pea06a, She07]. Cambridge [E.07g]. Canary [Gal09]. Canorus [Lat02, Lat05]. Capacities [Eld09]. Capacity [LP04]. Carabus [Kri08]. Case [Har09c, Pea06f, Per04b]. Cases [Bro07, Mac02b, Pea06h, Pea07b, Pea09c, Tur07]. Categories [Wel02b]. Cattle [BP06a]. Celandine [Ano03b]. Cell [War03]. Certain [Pea06a, Pea08a, Pea06f, Pea07c]. Change [Lut03, Pea02a, W.02, Wel01a]. Character [Pea02b, Pea06d, Pea09c]. Characteristics [Toc07]. Characters [JLP03, P.07a, PL03, Pea04b, Pea04a, Pea06f, Pea06i, Pea08d, SE07, W.02]. Charlier [E.06]. Chief [Rhi09]. Children [Eld09, Toc08, Toc09]. Chilomonas [Pea06c]. China [RP06]. Chrysanthemum [PY02, Tow02]. Cichorium [dH06]. Circa [Pea02a]. Circumferences [LP04]. Class [Pea06e]. Classification [Som06]. Classified [Toc06b]. Clausilia [Wel01b, Wel04]. Claw [Pea08c]. Co [Tow02]. Coat [BLP04, BP06a, Bla02]. Coat-Colour [BLP04, BP06a, Bla02]. Cobb [Pea08b]. Coefficient [Cob08, Har09c, Pea06e, Stu08a]. Cold [Dim02]. Collected [Sch06a].

Collection [Sch05a]. Collections [Pea03a]. Coloration [Lat02]. Colour [Ano04c, BLP04, BP06a, Bla02, P.04b, Pea09b, dC04]. Comb [PP09]. Common [Pea09d]. Communication [Jen09]. Comparative [Sch05a]. Comparison [Pea04b, TBP09]. Compound [P.07a]. Computation [Gib06, Rhi09]. Conceptions [Pea02d]. Concerning [Pea08a]. Conditions [Pea06c]. Coniferennadeln [Hey09, Hey09]. Conjugation [Pea07a]. Connected [Pea06f]. Constants [Ano03d, BLP05, Pea06a, Pea09a, Rhi09]. Contingency [BP06b, Per04b]. Contributions [Tsc05a]. Convenient [Yul04]. Cooperative [Ano02a, Ano03a, Ano03b, Ano06b, TBP09, WLP07]. Corpuscles [Pea09d]. Correcting [Ano04b]. Correction [Bat04, Bat05]. Correlation [Ano03b, Ano04c, Ano08c, Cob08, FL02, Gre04, Har09a, Har09c, JLP03, Lee02, LW02, Mac04, P.04b, PD03, Pea05a, PF05, Pea06f, PL08, Pea08d, Pea09c, Stu08a, Tsc05a, Tur06, Wel01a]. Correlationen [Ano08e]. Corresponding [Gal07]. Counting [Stu07]. Counts [GW09]. Crania [FL02, Mac04, Mac06, Mye03, Pea03a]. Cranial [LP04, vT03]. Craniological [Ano03c, JLP03, Mye03, P.03a, P.03b, Pea03a, Pea03b, vT03]. Criminal [Mac02a]. Criminals [GD04, Mac02a]. Criticisms [Pea07c]. Crosses [BP06a, Pea09b]. Crossing [Dar02, Dar03a, Dar03b, Dar04, Sch05b]. Cuckoo [Lat02, Lat05]. Cuculus [Lat02, Lat05]. Curve [E.07d, Pea06a, dH04, dH05]. Curves [Ano08b, Pea02f, Pea02e, Pea06g, Yul04, dH04, dH05]. Cynthia [Cra04]. Cytoplasm [Smi03]. Czekanowski [Ano08e, Ano08h].

D [War09]. Dangers [Per04a]. Daphnia [War03]. Darstellung [E.06]. Data [LW02, PP09, Pow01, Pow05, Sch06a, Toc06b]. Davenport [E.09a]. Deafness [Sch06a]. Death [Yul02a]. Deformities [Lew08a, LE08]. Deformity [Pea08c]. Department [Sch05a]. Dependent [Har09a]. Derselben [Sch08]. Describing [Pea06g]. Determination [Tow02]. Determining [Ano08g, Pea09c]. Development [Edw08]. Deviates [Gal07]. Deviation [Har09a, Pea08a]. Differences [Eld02a]. Differentiation [PR04]. Dimension [Eld08]. Dimensions [Lat02, Lat05]. Dimorfiche [dH04, dH05, E.07d]. dimorphic [dH04, dH05]. Discussion [Bro07, Sch06a]. Disease [Bro07]. Diseased [Gre04]. Distribution [Pea06h, Pea07b, She07, Stu09, Tur07]. Distributions [Bla05, GW09, Rhi09]. Divergence [Pea06e]. Divers [Ano03b]. Dolichocephalie [P.07b]. Domestic [PP09]. Domesticus [Pea02c]. Dory [Bry02]. Double [Pea08d]. Dr [Lee02, Sch06a]. Dr. [P.03a]. Drawing [Eld09, Yul04]. Drawn [Pea06d, Stu09]. Drones [WLP07]. Duration [BP01, Pea06b, Pow05]. Durch [Pea05b]. During [PY02, Wel01a, Pea02a]. Dying [Pea06b].

Each [Gal07, Lat05, Pea09c]. Earthworm [PF05]. ed [dH06]. Edgeworth [E.06]. Editorial [Ano01a, Ano01b, Ano02b, Ano03d]. Effect [Cob08, Dim02]. Egg [Jen06, Jen09, Lat02, Lat05, Pea02c]. Eggs [Lat02]. Eichorni [Smi03]. Elementary [Ano04b]. Elimination [Cra04]. Embryo [Jen06, Jen09]. English [BLP05, Mac04, Mac06]. Enquiry [HM09, Lat02]. Entrants [E.09b]. Environment [Dim02]. Epidemic [Tur06]. Ernest [Pea09d]. Error [BP06b, E.06, Pea08a, PL08, Stu07, Stu08a, Stu08b, E.06]. Errors [Ano03d, Cob08, Gib06, Pea06a, Rhi09]. Estimating [Bro07]. Etc [Ano02b]. Étude [Gai04]. Eugene [E.09b]. Eugenics [E.08, E.07e]. Eupagurus [Sch03]. European [Dar02, Dar03a]. Evolution [LW02, Pow01, Pow05, E.07f]. Exceeds [Pea09c]. Exhibiting [Lat05]. Expectation [Pea02a, E.07c]. Experience [E.07c]. Experimental [Cra04]. Explanatory [Ano06c]. External [BLP05]. Extracted [Dar03b]. Extrapolation [Per04a]. Extremes [E.09b]. Eye [P.04b]. Eyes [Ano04c].

F [E.06, McI03, Wel07]. Facilitate [Rhi09]. Facilitating [Gib06]. Fagus [PR04]. Falls [Pea09c]. Fat [Rie09]. Fathers [Per04b]. Favourable [Pea06c]. Fay [Sch06a]. Fechner [Pea05b]. Fecundity [Sur09]. feet [Sch08]. Fehlergesetz [Pea05b, E.06]. Fehlergesetzes [E.06]. Female [RP06]. femmes [Gor09]. Fernando [E.07d]. Fertility [Dav02, Pow05]. Ficaria [Wel01a]. Figure [Gal07]. Finger [Ano03e]. Finite [Eld02a]. First [Di 07, E.07a, Gre04, GW09, S02, Wel01b]. Fisk [E.09b]. Fission [Sim02]. Fit [Eld02b]. Fitting [Pea02f, Pea02e]. Fleur [Gai04]. Flinders [E.07b]. florescences [dH06]. Floridana [Edw08]. Flower [Rob08, Gai04]. Flower-heads [Rob08]. Flowering [Wel01a]. Flowers [PY02, Tow02]. Foot [Lew08a, LE08, Pea08c]. Force [E.09b]. Form [E.06]. Forms [Dim02]. Formula [Ano08g]. Formulae [Ano04b]. Foster [Lat02, Lat05]. Foster-Parent [Lat02, Lat05]. Foundation [E.07e]. Fourteenth [E.08]. Fowl [PP09]. Frage [Ano08e, Sch08]. Francis [E.07e, Ano02d, Ano02e]. Frank [Ano06i]. Frederick [P.09]. French [Gai04]. Frequency [Ano03d, Bla05, GW09, Pea09a, Pea06g, Rhi09, She07, Som06, Ano08b]. Frequency-Distribution [She07]. Frequency-Ratios [Som06]. Frog [Jen06, Jen09]. Front [Ano02c, Ano04a, Ano05a, Ano06e, Ano06d, Ano08f]. Function [Duf09]. Functionen [E.06]. Fundamental [Pea02d]. Further [Bat05, Mac03, Pea07b]. Füsse [Sch08]. Future [E.07c].

G [Ano08c, P.07a, Pea07c, War09]. G. [PY02]. Gaillardia [Rob08]. Galton [Ano02d, Ano02e, E.07e, Sch06b]. Gathered [PY02]. Gelasimus [Dun03, Dun03]. General [TBP09]. Generalised [E.06, PL08]. Generations [War02]. German [Ano08h, Dun03, Hen02, Hey09, Lud01, Sch08]. give [Pea06f]. Given [Pea09c]. Goodness [Eld02b]. Grade [Gal07, Pea09c]. Grades [Gal07]. Graduation [Eld03a, Eld03b, Spe04]. Grandchildren [Bla02]. Grandsire [Bla02]. Great [E.06]. Greene [Tow02]. Grey [Sch05b]. Greyhound [BLP04]. Growth [Edw08].

H [E.07f]. Haemacytometer [Stu07]. Hair [Ano04c, P.04b]. Hand [LW02, Lew08a, LE08]. Harbor [Dim02]. Haviland [War09]. Head [Gla05, Pea06i, Ano08h]. Heads [PY02, Rob08]. Healthy [Gre04]. Heart [Gre04]. Heinrich [Ano09b]. Helguero [E.07d]. Helicogena [Di 07]. Helix [Hen02, Di 07, Hen02]. Heller [Sch03]. Henderson [Ano08b]. Hereditary [Sch06a]. Heredity [E.07f, Lut03, P.09, PYBL03, Pea08b, Wel03, Wel07]. Heritance [Ano06b]. Heterogeneity [Mye03, Pea03a, Pea06e]. Hétérostylie [Gai04]. heterostyly [Gai04]. History [P.09]. Hoc [Ano01c]. Holothuria [Edw08]. homicides [Gor09]. Homogeneity [Mye03, Pea03a]. Homotyposis [PR04]. Horses [Bla02]. House [Sch05b]. Human [FL02, Gre04, Mac04]. Hundred [GD04]. Huxley [E.07b]. Hyalopterus [War02]. Hybrids [Dar03b]. Hypothesis [Spe04].

Identification [Mac02a]. if [Lat05]. Ignoramus [Ano01c]. II [Ano03b, Pea02f, Pea04b]. III [Ano06b]. Illustration [Ano02d]. Immunity [Bro05, Bro07]. Importance [E.08]. Including [Gal07]. Index [Ano09c, GW09, HM09, P.07b]. Index-Angaben [P.07b]. Individuals [Pea06b]. Inflorescenze [dH06]. Influence [E.07c, Lut03, Mac02b, Pea08d]. Information [Ano01d, Ano02f, Ano04d, Ano05b, Ano06h, Ano06g, Ano08i, Ano09f]. Inheritance [Ano02a, Ano03e, BLP04, BP06a, BP01, Bla02, Her06, Her08, Lew08b, P.04a, P.07a, P.08, PL03, Pea04b, Pea04a, Pea08c, Rie09, RP06, SE07, W.03, War02, Wel02a, Wel07, Woo06, PL03, Sch06b]. Inmates [Toc07]. Inoculation [May09]. Integral [Har09c, She03]. Intelligence [Pea06i]. Intensity [BP01, Lut03, Pea09c]. Internal [LP04]. Interpolation [Eld02a, Eld08]. Interracial [JLP03, Tsc05a]. Intybus [dH06]. Investigations [Ano02a, Ano03b, Ano06b]. Island [Dim02]. Itala [Wel04]. Italian [dH04, dH05, dH06].

J [Ano08d]. Jager [Edw08]. Jan [Ano08e, Ano08h]. Japanese [Dar02, Dar03a, Dar03b, Dar04]. Johannsen [P.07b]. John [Bry02, E.07g]. July [PY02]. June [E.07g].

K. [PY02]. Kapteyn [Pea06j]. Karl [E.07a, E.07c, E.08]. Keswick [PY02]. Key [Ano06f]. Known [Pea08c]. Kopfmasse [Ano08h]. Kritik [P.07b].

L [E.06, E.07g, E.09b, Kri08]. L. [Gai04, Tow02, dH06]. Laboremus [Ano01c]. Laminata [Wel01b]. Langenvariation [Hey09]. Late [War09, Wel07]. Latr [Dun03, Dun03]. Law [E.06, P.04a, PYBL03, Sch06b]. Laws [PL03, Pea04b, W.04, Wel02a, Woo03, P.07a]. Leaves [Fry02, PR04]. Lecture [E.07b, E.08]. Length [Gal07, Hey09]. Lepidoptera [Cra04]. Lesser [Ano03b]. Leucanthemum [Tow02, PY02]. Leucosis [Ano09d]. Life [BP01, Pea06b, Pea02a, Pow05, E.09b]. Line [Pea06f, RP06]. Linear [Pea06f]. Linearity [Bla05]. Litter [RP06, Wel07]. Lobster [Pea08c]. Lobster-Claw [Pea08c]. Local [Yul02a]. Localities [Ano03b]. Lock [E.07f]. London [Tur06]. Long [Dim02, Sch05a]. Ludwig [Lee02].

M [E.07b]. Magna [War03]. Magnifica [P.01]. Mainly [War09]. Makeham [Spe04]. Man [Ano03a, Ano04c, BP01, JLP03, LW02, P.04b, P.08, Pea05a, Pea02a, PL03, Pea04b, Pow01, Pow05, Tsc05b, Woo06]. Mantis [dC04]. Map [Ano06f]. Martens [Wel04]. Mass [Smi03, Ano08h, Ano08h]. Massimi [dH04]. Materialien [Lud01]. materials [Lud01]. Mating [Ano03a]. Matter [Ano02c, Ano04a, Ano05a, Ano06e, Ano06d, Ano08f]. maxima [dH04]. Mean [BP06b, Pea06b, Pea03b, Stu08b]. Mean-Square [BP06b]. Means [Bro07, Stu09, Yul04]. Measured [Pea09c]. Measurement [LP04]. Measurements [BLP05, GD04, Pea02e, Pea09d, Toc06a]. Measurments [Pea02f]. Medico [HM09]. Medico-Statistical [HM09]. Memoir [Ano06b, GW09, Lat05, Lew08a]. Mendel [P.07a, W.04, Wel02b, Wel02a, Wel03, Woo03]. Mendelian [P.04a]. Mental [Pea04b, Pea06i, P.09]. Merism [Pun04]. Meristic [Pea04a]. Method [Har09c, Pea09c]. Mice [Dar02, Dar03a, Dar03b, Dar04, Sch05b, Wel07]. Migrations [E.07b]. Millesimal [Gal07]. Modal [Pea02b]. Modes [Tow02]. Moments [Ano04b, She07, Ano08b]. Montagu [Wel01b]. Moorfields [Mac06]. Moral [P.09, Pea04b]. Moscatel [Whi02]. Most [Pea06g, S02]. Mr [Wel03, Pea04a, Pea07c, Pea08b, War09]. Muhl [W.02]. Mulberry [Fry02]. Muller [McI03]. Multiple [PL08]. Museum [Sch05a]. Muskelvarietaten [Ano08e]. Myers [P.03a]. Myres [E.07g].

Naqada [FL02]. Nassa [Dim02]. National [E.08]. Natural [BP01, Di 07, Eld03c, Wel01b]. Negro [Pea09b]. nelle [dH06]. Nemoralis [Hen02, Hen02]. Nemorosa [Yul02b]. Nest [TBP09]. Nichols [Ano08d]. Niger [Pun04]. Nigra [McI03]. No [E.07g]. Non [Pea06d, Woo06]. Non-Inheritance [Woo06]. Non-Significant [Pea06d]. Normal [Gal07, Pea06a, PL08]. Notation [Ano08c]. Note [Ano06c, Ano09d, Dar02, Fry02, Har09b, JLP03, Kri08, Lew08b, Lut03, P.03a, P.03b, P.08, Pea04a, Pea06d, Pea08b, Pea09b, Wel04, dC04, vT03]. Notes [Ano03c, Eld05, Eld08, JLP03, Lat02, Mye03, P.03a, P.03b, Pea03a, Pea03b, Yul03, vT03]. Nucleus [Smi03]. Number [Ano08c, Bry02, Yul02b]. Numbers [Eld03c, E.06]. Numerical [Ano08d].

O [McI03]. O. [Tow02]. Observation [Cob08, Eld02b]. Observations [Edg07, Pea02f, Pea02e, War09]. Obsoleta [Dim02]. Occupation [Per04b]. Officinalis [Gai04]. Omotiposi [dH06]. One [GD04, Lat05]. Only [Pea09c]. Ophiocoma [McI03]. Opsonic [GW09, HM09]. Order [Pea06f]. Organ [Pea02b]. Organic [Wel01a]. Origin [LE08, Wel02c]. Other [Ano04b, Eld09, Pea06i]. Oxford [Sch05a, E.07g].

Paramecium [Pea07a]. Parent [Lat02, Lat05]. Part [BP06a, Pea02f]. Parthenogenetic [War02]. Partial [Ano09d, Gal09]. Passer [Pea02c]. Past [E.07c]. Pearson [E.07a, E.07c, E.08, PY02]. Peas [Wel02a]. Pectinatella [P.01]. Pendulations [Ano09b]. Pendulations-Theorie [Ano09b].

Percentage [Pea09c]. Period [Pea02a]. Petrie [E.07b]. Phagocytic [GW09]. Phagocytosis [GW09]. Phaseolus [W.03]. Philosamia [Cra04]. Photograph [Ano01c, Ano02e]. Physical [BLP05, PL03, Pea04b, Pea06i]. Pigmentation [Toc08, Toc09]. Plants [Ano02a, Ano02b, Ano03b, Ano06b, Lee02]. Plates [Bry02]. Points [Ano04b, Pea08a, Pea06f]. Poland [RP06]. Polymorphism [Ano02b]. Poppies [Ano09e]. Poppy [Ano02a, Ano06b]. Population [Pea06g, TBP09]. Pourtales [Edw08]. Powers [Eld03c]. Preference [Lat05]. Preliminary [JLP03, War03, dC04]. Prenanthoides [W.02]. Pretoria [Ano09e]. Previous [Mac02b]. Prideauxi [Sch03]. Principles [E.09a]. Prints [Ano03e]. Prizes [S02]. Probability [She03, E.06, E.07e]. Probable [Ano03d, BP06b, Gib06, Har09a, Pea06a, Pea08a, PL08, Rhi09, Stu08a, Stu08b]. Problem [LW02, Pow01, Pow05]. Probleme [Lud01]. problems [Lud01]. Processes [Eld05]. Production [Rie09]. Prof [E.07b]. Professor [P.03b, Pea03b, Pea06j, Sch06b, Wel02c]. Progress [E.07f]. Proof [Ano04b]. Proportion [S02]. Proportions [Ano08d]. Protective [dC04]. Psychical [SE07]. Psychology [P.09]. Pugilator [Dun03]. Pulmonaria [Gai04]. Pulmonary [E.07a]. Punnett [Pea04a]. Pure [BP06a].

quadrapartite [Sch08]. Quantitative [Dim02, Tsc05b]. Quantitatively [P.07a]. Queens [TBP09, WLP07]. question [Sch08].

R [E.07f, Wel07]. Rabbit [Woo03]. Race [Wel04]. Racehorse [Her08]. Races [Dar02, Dar03a, Dar03b]. Rana [Jen06, Pea09d]. Random [Pea06g, Stu09]. ranunculoides [Wel01a]. Raphael [Ano06i]. Rates [Yul02a]. Ratio [Her06, Her08, Wel07, Ano08h]. Ratios [Som06]. Raw [Ano04b]. Ray [PY02, Tow02]. Ray-Flowers [PY02, Tow02]. Recessive [Dar03b]. Recorded [Pea09c]. Records [Gal09, Wel07, Woo03]. Red [Pea09d]. Reference [FL02, Gre04, GW09, Mac04, Mac06]. Regeneration [Sch08]. Regression [Bla05, Pea06f]. Rejoinder [Pea05b, Pea06j]. Relating [Mac03]. Relation [Jen06, Jen09, Lat02, Sim02]. Relations [Gla05, Smi03]. Relationship [Pea06i, War03]. Relationships [BLP05]. Relative [Kri08]. Religiosa [dC04]. Remarks [P.03a, P.03b, Pea07b, Tow02, Pea08b]. Reply [Pea07c]. Report [Dar03a, Dar03b]. Reproductive [Pow05]. Researches [E.06]. Resemblance [Tsc05b]. resolution [dH05]. Result [Dar03a, Dar03b, Dar04]. Results [Dar02, Sch05b]. Review [Ano08e, Ano08d, Ano08c, Ano08b, Ano08h, Ano09b, E.07e, E.07a, E.07c, E.07f, E.07b, E.07d, E.08, E.09b, E.09a, Gor09, P.07a, P.07b, P.09, Ano08h, E.07g]. Reviews [E.06]. Revisions [Wel03]. Risoluzione [E.07d, dH05]. Robert [Ano08b]. Round [Sch05a]. Royalty [P.09].

S [P.03a]. same [Sch08]. Sample [Pea06d]. Samples [Pea06g, Stu09]. Scale [Pea06f]. Scales [Yul04]. Schadelmassen [Ano08h]. School [Eld09, Toc08, Toc09]. Science [E.08]. Scope [Ano01a, E.08]. Scotland [Toc07, Toc08, Toc09]. Scottish [Ano06f]. Season [Wel01a]. Seasonal [W.02]. Second [Ano06b, Dar03a, FL02, Lat05, LW02, Mac06, S02, TBP09, Jen09]. Section [Pea04a]. Segmentation [Jen09]. Seine [Pea05b]. Selection [BP01, Di 07, E.09b, Pea08d, Pow05, Wel01b]. Sepals [Yul02b]. Severity [Bro07, Pea06h, Pea07b, Tur07]. Sex [Her06, Her08, Lut03, Pea08b, Pun04, Wel07, Woo06]. Sex-Ratio [Her06, Her08, Wel07]. Sexes [Ano08d, Kri08]. Shape [Pea06i]. Sheppard [Ano04b]. Shirley [Ano02a, Ano06b, Ano09e]. Short [Har09c, Pea09c]. Shorthorn [BP06a, BP06a]. Shorthorns [BP06a]. Sicily [Bat04, Bat05, W.04]. Sickness [Eld03a, Eld03b, Spe04]. Significant [Pea06d]. Signo [Ano01c]. Simroth [Ano09b]. Single [TBP09]. Size [Gla05, Lat02, Pea06i, RP06, War03, Wel07]. Skew [Rhi09]. Skin [Pea09b]. Skin-Colour [Pea09b]. Skull [FL02, Mac04, Mac06, Ano08h]. Skulls [Sch05a]. Smallpox [Mac03, Mac02b, Pea06h, Pea07b, Tur06, Tur07, Bro05]. Some [Eld08, Lat05, Pea06f, Per04a, War09, Woo03]. Sons [Per04b]. Sources [Ano02b]. Sows [RP06]. Space [Eld08]. Sparrow [Pea02c]. Special [FL02, Gre04, GW09, Mac04, Mac06]. Species [Dav02, Lat05, Wel02c]. Spielarten [Hen02]. Spinax [Pun04]. Spirit [Ano01b]. Split [Lew08a, LE08, Pea08c]. Split-Foot [Lew08a, LE08, Pea08c]. Split-Hand [Lew08a, LE08]. Spring [Dim02]. Springs [Tow02]. Square [BP06b]. Standard [Pea08a]. State [E.08]. Statistical [Bro05, Bro07, Cra04, Edg07, Eld05, HM09, May09, P.09, War09, Lud01]. Statistics [Mac03, Yul03]. Statoblasts [P.01]. Straus [War03]. Studies [Bro05, Bro07, Cra04, Pea05a, Hen02, Ano08h]. Study [Ano03a, BLP05, Di 07, Dim02, E.07a, E.07f, FL02, Gla05, Gre04, GW09, LW02, Mac03, Mac04, Mac06, May09, P.09, Pea07a, Pea09d, Smi03, TBP09, Tsc05a, Tsc05b, Wel01b, WLP07, Gai04]. Sub [Lat05, Pea06d]. Sub-Sample [Pea06d]. Sub-Species [Lat05]. Sui [dH04]. Suitable [Pea06g, S02]. Sums [Eld03c]. Survey [Toc08, Toc09]. Swine [Sur09]. Sylvatica [PR04]. Symmetry [Dav02, Jen06, Jen09]. System [Ano08c]. Systematic [Pea02f, Pea02e].

Table [Eld03a, Eld03b, Gal07, Spe04]. Tables [Ano06c, Duf09, Eld02b, Eld03c, Gib06, Rhi09, She03, Toc06b]. Tadpole [Pea09d]. Tarsus [Sch08]. Temporaria [Jen06, Pea09d]. Tending [Lat05]. Termites [War09]. Testing [Eld02b]. Tests [Bla05]. Their [BLP05, LE08, JLP03]. Theorie [Ano09b]. Theory [Ano08c, E.06, Eld02b, P.07a, Wel03, Yul03]. Third [Dar03b]. Thirty [GD04]. Thoroughbred [Bla02, Her08]. Torok [P.03b, Pea03b]. Transmission [LE08]. treated [Ano08c]. Trees [Fry02]. Trihodus [War02]. Trivittata [Dim02]. Tuberculosis [E.07a]. Two [Pea06f, Pea08d]. Types [LE08, vT03]. Typhoid [May09].

U [PY02, Pea07c]. Udny [Ano08c, P.07a]. Unfavorable [Pea06c]. Untersuchungen [Hen02, Ano08h]. Upon [Cob08, Cra04, Dim02, Tow02].

V [E.06, Ano09c]. Vaccination [Bro05, Mac03, Mac02b, Tur06]. Value [Har09a, Pea02b, S02, dC04]. Variability [Ano09e, Dav02, Gre04, Kri08, LW02]. Variable [Har09a, Pea09a]. Variables [Ano08c]. Variates [Har09c]. Variation [Bro01, E.07f, Edw08, FL02, Fry02, Har09b, Lee02, Mac04, McI03, PD03, PF05, Pea06c, PP09, Pea02c, PY02, Pea08d, Rob08, Sch03, Sim02, Tow02, War02, Whi02, Yul02b, Hey09, Ano03b, Cra04, Pea05a, dH06]. Variational [Lud01]. Variational-statistical [Lud01]. Variations [Gai04, Lat02]. Variationsstatistische [Lud01]. Variazione [dH06]. varieties [Hen02]. Various [Yul04]. Verallgemeinerungen [Pea05b]. Verhaltnis [Ano08h]. Vespa [TBP09, WLP07]. Viergliederigen [Sch08]. View [P.04a]. Viscera [Gre04]. Volume [Ano01d, Ano02f, Ano04d, Ano05b, Ano06h, Ano06g, Ano08i, Ano09f]. Volumes [Ano09c]. Vom [Sch08]. Vries [Wel02c]. Vulgaris [W.03, WLP07, TBP09].

W [E.07b, P.07b]. W. [Wel07]. Walker [War02]. Walter [Ano06i]. Waltizing [Dar02]. Waltzing [Dar03a, Dar03b, Dar04]. Warren [Pea09d]. Wasps [Edg07]. Weight [Gre04, Pea05a]. Weights [BLP05]. Weldon [Ano06i, Wel07]. Wherein [Pea09c]. which [Pea06f, Pea06g, Pea09c, Stu09]. White [Pea09b]. Willkurlicher [E.06]. within [Pea06b]. Woman [Tsc05b]. Woods [P.09]. Work [War09]. Workers [WLP07].

Year [Pea06b]. Years [Pea02a]. Yellow [Tow02]. Young [Bry02]. Yule [Ano08c, P.07a, PY02, Pea07c].

Ziegler [Sch06b]. Zur [P.07b, Sch08, Ano08e]. Zweite [E.06].

References

Anonymous:1901:ESBa

[Ano01a] Anonymous. Editorial: The scope of Biometrika. Biometrika, 1(1): 1–2, October 1901. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

        3510 (electronic). URL http://www.jstor.org/stable/2331667.

Anonymous:1901:ESBb

[Ano01b] Anonymous. Editorial: The spirit of Biometrika. Biometrika, 1(1): 3–6, October 1901. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

        3510 (electronic). URL http://www.jstor.org/stable/2331668.

Anonymous:1901:PIH

[Ano01c] Anonymous. [photograph]: Ignoramus, in hoc signo laboremus. Biometrika, 1(1):??, October 1901. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331666.

Anonymous:1901:VI

[Ano01d] Anonymous. Volume information. Biometrika, 1(1):i–iv, October 1901. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331665.

Anonymous:1902:CIP

[Ano02a] Anonymous. Cooperative investigations on plants: I. On inheritance in the Shirley poppy. Biometrika, 2(1):56–100, November 1902. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331501.

Anonymous:1902:ESA

[Ano02b] Anonymous. Editorial: On the sources of apparent polymorphism in plants, etc. Biometrika, 1(3):304–306, April 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331541.

Anonymous:1902:FM

[Ano02c] Anonymous. Front matter. Biometrika, 2(1):i, November 1902. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331495.

Anonymous:1902:IFG

[Ano02d] Anonymous. [illustration]: Francis Galton. Biometrika, 2(1):??, November 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331497.

Anonymous:1902:PFG

[Ano02e] Anonymous. [photograph]: Francis Galton. Biometrika, 2(1):??, November 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331496.

Anonymous:1902:VI

[Ano02f] Anonymous. Volume information. Biometrika, 2(1):ii–iv, November 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331494.

Anonymous:1903:AMM

[Ano03a] Anonymous. Assortative mating in man: A cooperative study. Biometrika, 2(4):481–498, November 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331510.

Anonymous:1903:CIP

[Ano03b] Anonymous. Cooperative investigations on plants: II. Variation and correlation in lesser celandine from divers localities. Biometrika, 2(2):145–164, February 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331679.

Anonymous:1903:CN

[Ano03c] Anonymous. Craniological notes. Biometrika, 2(3):338–339, June 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331609.

Anonymous:1903:EPE

[Ano03d] Anonymous. Editorial: On the probable errors of frequency constants. Biometrika, 2(3):273–281, June 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331603.

Anonymous:1903:IFP

[Ano03e] Anonymous. Inheritance of finger prints. Biometrika, 2(3):356, June 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331613.

Anonymous:1904:FM

[Ano04a] Anonymous. Front matter. Biometrika, 3(1):ii, January 1904. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331518.

Anonymous:1904:EPS

[Ano04b] Anonymous. On an elementary proof of Sheppard’s formulae for correcting raw moments and on other allied points. Biometrika, 3 (2/3):308–312, March/July 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331483.

Anonymous:1904:CBA

[Ano04c] Anonymous. On the correlation between age and the colour of hair and eyes in man. Biometrika, 3(4):462–466, November 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331732.

Anonymous:1904:VI

[Ano04d] Anonymous. Volume information. Biometrika, 3(1):i–vi, January 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331517.

Anonymous:1905:FM

[Ano05a] Anonymous. Front matter. Biometrika, 4(1/2):ii, June 1905. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331530.

Anonymous:1905:VI

[Ano05b] Anonymous. Volume information. Biometrika, 4(1/2):i–iv, June 1905. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331529.

Anonymous:1906:B

[Ano06a] Anonymous. Bibliography. Biometrika, 5(1/2):210–212, October 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331664.

Anonymous:1906:CIP

[Ano06b] Anonymous. Cooperative investigations on plants: III. On heritance in the Shirley poppy: Second memoir. Biometrika, 4(4):394–426, March 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331637.

Anonymous:1906:ENT

[Ano06c] Anonymous. Explanatory note to tables. Biometrika, 5(supplement): 1, ???? 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331577. Anthro-

pometric survey of the inmates of the asylums in Scotland.

Anonymous:1906:FMb

[Ano06d] Anonymous. Front matter. Biometrika, 5(supplement):??, ???? 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331576. Anthropometric sur-

vey of the inmates of the asylums in Scotland.

Anonymous:1906:FMa

[Ano06e] Anonymous. Front matter. Biometrika, 5(1/2):ii, October 1906. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331644.

Anonymous:1906:MSA

[Ano06f] Anonymous. [map of Scottish asylums and key to map]. Biometrika, 5(supplement):2–3, ???? 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2331578. Anthropometric survey of the inmates of the asylums in Scotland.

Anonymous:1906:VIb

[Ano06g] Anonymous. Volume information. Biometrika, 5(supplement):i–xviii, ???? 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331575. Anthro-

pometric survey of the inmates of the asylums in Scotland.

Anonymous:1906:VIa

[Ano06h] Anonymous. Volume information. Biometrika, 5(1/2):i–vi, October 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331643.

Anonymous:1906:WFR

[Ano06i] Anonymous. Walter Frank Raphael Weldon. 1860–1906. Biometrika, 5(1/2):1–52, October 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331645.

Anonymous:1907:B

[Ano07] Anonymous. Bibliography. Biometrika, 5(4):482–486, June 1907. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331703.

Anonymous:1908:B

[Ano08a] Anonymous. Bibliography. Biometrika, 6(1):124–128, March 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331574.

Anonymous:1908:BRBd

[Ano08b] Anonymous. Book review: Frequency Curves and Moments. by Robert Henderson. Biometrika, 6(1):124, March 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331572.

Anonymous:1908:BRBc

[Ano08c] Anonymous. Book review: On the Theory of Correlation for any Number of Variables treated by a New System of Notation, by G. Udny Yule. Biometrika, 6(1):124, March 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331571.

Anonymous:1908:BRBb

[Ano08d] Anonymous. Book review: The Numerical Proportions of the Sexes at Birth, by J. B. Nichols. Biometrika, 6(1):123, March 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331570.

Anonymous:1908:BRBa

[Ano08e] Anonymous. Book review: Zur Frage der Correlationen der Muskelvarietaten. by Jan Czekanowski. Biometrika, 6(1):123, March 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331569.

Anonymous:1908:FM

[Ano08f] Anonymous. Front matter. Biometrika, 6(1):ii, March 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331553.

Anonymous:1908:FD

[Ano08g] Anonymous. On a formula for determining Γ(x + 1). Biometrika, 6(1):118–119, March 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331566.

Anonymous:1908:RBV

[Ano08h] Anonymous. Review: Untersuchungen über das Verhaltnis der Kopfmasse zu den Schadelmassen by Jan Czekanowski. (German) [Review: Studies on the ratio of the head mass to the skull mass, by Jan Czekanowski]. Biometrika, 6(1):123, March 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331568.

Anonymous:1908:VI

[Ano08i] Anonymous. Volume information. Biometrika, 6(1):i–vi, March 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331552.

Anonymous:1909:B

[Ano09a] Anonymous. Bibliography. Biometrika, 6(4):445–451, March 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331469.

Anonymous:1909:BRB

[Ano09b] Anonymous. Book review: Die Pendulations-Theorie, by Heinrich Simroth. Biometrika, 6(4):444, March 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331467.

Anonymous:1909:IVV

[Ano09c] Anonymous. Index to Volumes I–V of Biometrika. Biometrika, 6(4): i–xviii, March 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331471.

Anonymous:1909:NPL

[Ano09d] Anonymous. Note on partial leucosis in a hen. Biometrika, 7 (1/2):234–236, July/October 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2345375.

Anonymous:1909:VSP

[Ano09e] Anonymous. Variability in Shirley poppies from Pretoria. Biometrika, 7(1/2):227–230, July/October 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2345373.

Anonymous:1909:VI

[Ano09f] Anonymous. Volume information. Biometrika, 7(1/2):i–v, July/ October 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2345360.

Bateson:1904:ASC

[Bat04] W. Bateson. Albinism in Sicily — a correction. Biometrika, 3 (4):471–472, November 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331735. See [W.04].

Bateson:1905:ASF

[Bat05] W. Bateson. Albinism in Sicily. A further correction. Biometrika, 4(1/2):231–232, June 1905. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331539. See [W.04].

Blanchard:1902:ICC

[Bla02] N. Blanchard. On the inheritance in coat-colour of thoroughbred horses (grandsire and grandchildren). Biometrika, 1(3):361–364, April 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331548.

Blakeman:1905:TLR

[Bla05] John Blakeman. On tests for linearity of regression in frequency distributions. Biometrika, 4(3):332–350, November 1905. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331617.

Barrington:1904:ICC

[BLP04] Amy Barrington, Alice Lee, and Karl Pearson. On inheritance of coat-colour in the greyhound. Biometrika, 3(2/3):245–298, March/ July 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331481.

Blakeman:1905:SBC

[BLP05] J. Blakeman, Alice Lee, and Karl Pearson. A study of the biometric constants of English brain-weights, and their relationships to external physical measurements. Biometrika, 4(1/2):124–160, June 1905. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331534.

Beeton:1901:IDL

[BP01] Mary Beeton and Karl Pearson. On the inheritance of the duration of life, and on the intensity of natural selection in man. Biometrika, 1(1):50–89, October 1901. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331672.

Barrington:1906:ICC

[BP06a] A. Barrington and K. Pearson. On the inheritance of coat-colour in cattle: Part I. Shorthorn crosses and pure shorthorn crosses and pure shorthorns. Biometrika, 4(4):427–464, March 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331638.

Blakeman:1906:PEM

[BP06b] John Blakeman and Karl Pearson. On the probable error of meansquare contingency. Biometrika, 5(1/2):191–197, October 1906. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331660.

Browne:1901:VBA

[Bro01] Edward T. Browne. Variation in Aurelia Aurita. Biometrika, 1(1):90– 108, October 1901. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331673.

Brownlee:1905:SSI

[Bro05] John Brownlee. Statistical studies in immunity. Smallpox and vaccination. Biometrika, 4(3):313–331, November 1905. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331616.

Brownlee:1907:SSI

[Bro07] John Brownlee. Statistical studies in immunity: A discussion of the means of estimating the severity of cases of acute disease. Biometrika, 5(4):423–435, June 1907. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331689.

Bryne:1902:NAB

[Bry02] L. W. Bryne. On the number and arrangement of the bony plates of the young John Dory. Biometrika, 2(1):115–120, November 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331506.

Cobb:1908:EEO

[Cob08] J. A. Cobb. The effect of errors of observation upon the correlation coefficient. Biometrika, 6(1):109, March 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331561.

Crampton:1904:ESS

[Cra04] Henry Edward Crampton. Experimental and statistical studies upon Lepidoptera: I. Variation and elimination in Philosamia Cynthia.

Biometrika, 3(2/3):113–130, March/July 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331478.

Darbishire:1902:NRC

[Dar02] A. D. Darbishire. Note on the results of crossing Japanese waltizing mice with European albino races. Biometrika, 2(1):101–104, November 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331502.

Darbishire:1903:SRR

[Dar03a] A. D. Darbishire. Second report on the result of crossing Japanese waltzing mice with European albino races. Biometrika, 2(2):165–173, February 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331680.

Darbishire:1903:TRH

[Dar03b] A. D. Darbishire. Third report on hybrids between waltzing mice and albino races: On the result of crossing Japanese waltzing mice with “extracted” recessive albinos. Biometrika, 2(3):282–285, June 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331604.

Darbishire:1904:RCJ

[Dar04] A. D. Darbishire. On the result of crossing Japanese waltzing with albino mice. Biometrika, 3(1):1–51, January 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331519.

Davenport:1902:VSF

[Dav02] C. B. Davenport. Variability, symmetry and fertility in an abnormal species. Biometrika, 1(2):255–256, January 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331489.

diCesnola:1904:PNP

[dC04] A. P. di Cesnola. Preliminary note on the protective value of colour in mantis religiosa. Biometrika, 3(1):58–59, January 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331521.

deHelguero:1904:SMD

[dH04] Fernando de Helguero. Sui massimi delle curve dimorfiche. (Italian) [On maxima of dimorphic curves]. Biometrika, 3(1):84–98, January 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331524.

deHelguero:1905:PRD

[dH05] Fernando de Helguero. Per la risoluzione delle curve dimorfiche. (Italian) [On the resolution of dimorphic curves]. Biometrika, 4(1/2):230– 231, June 1905. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331538.

deHelguero:1906:VEO

[dH06] Fernando de Helguero. Variazione ed omotiposi nelle inflorescenze di Cichorium Intybus L.. (Italian) [Variation and omotiposi (??) in florescences of Cichorium Intybus L.]. Biometrika, 5(1/2):184–189, October 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331658.

DiCesnola:1907:FSN

[Di 07] A. P. Di Cesnola. A first study of natural selection in “Helix Arbustorum” (Helicogena). Biometrika, 5(4):387–399, June 1907. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331686.

Dimon:1902:QSE

[Dim02] Abigail Camp Dimon. Quantitative study of the effect of environment upon the forms of Nassa Obsoleta and Nassa Trivittata from Cold Spring Harbor, Long Island. Biometrika, 2(1):24–43, November 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331499.

Duffell:1909:TF

[Duf09] J. H. Duffell. Tables of the Γ-function. Biometrika, 7(1/2):43–47, July/ October 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2345362.

Duncker:1903:ABP

[Dun03] Georg Duncker. Über Asymmetrie Bei “Gelasimus Pugilator” Latr. (German) [On antisymmetry in Gelasimus Pugilator” Latr].

Biometrika, 2(3):307–320, June 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331607.

E:1906:BRB

[E.06] W. P. E. Book reviews: The Law of Error by F. Y. Edgeworth; The Generalised Law of Error, or Law of Great Numbers; Über das Fehlergesetz by C. V. L. Charlier; Die Zweite Form des Fehlergesetzes; Über die Darstellung Willkurlicher Functionen; Researches into the Theory of Probability. Biometrika, 5(1/2):206–210, October 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331663.

E:1907:BRBb

[E.07a] W. P. E. Book review: A First Study of Pulmonary Tuberculosis, by Karl Pearson. Biometrika, 5(4):478, June 1907. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331696.

E:1907:BRBe

[E.07b] W. P. E. Book review: Migrations (The Huxley Lecture for 1906), by Prof. W. M. Flinders Petrie. Biometrika, 5(4):480, June 1907. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331699.

E:1907:BRBc

[E.07c] W. P. E. Book review: On the Influence of Past Experience on Future Expectation, by Karl Pearson. Biometrika, 5(4):478, June 1907. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331697.

E:1907:BRBf

[E.07d] W. P. E. Book review: Per la Risoluzione delle Curve Dimorfiche. by Fernando de Helguero. Biometrika, 5(4):480, June 1907. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331700.

E:1907:BRBa

[E.07e] W. P. E. Book review: Probability, the Foundation of Eugenics. by Francis Galton. Biometrika, 5(4):477, June 1907. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331694.

E:1907:BRBd

[E.07f] W. P. E. Book review: Recent Progress in the Study of Variation, Heredity and Evolution, by R. H. Lock. Biometrika, 5(4):478–480, June 1907. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331698.

E:1907:BCR

[E.07g] W. P. E. The Oxford and Cambridge Review, No. 1, June 1907. Article on “A Bureau of Biometry”, by John L. Myres. Biometrika, 5(4):477, June 1907. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331695.

E:1908:BRB

[E.08] W. P. E. Book review: The Scope and Importance to the State of the Science of National Eugenics. The Fourteenth Boyle Lecture, by Karl Pearson. Biometrika, 6(1):124, March 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331573.

E:1909:BRBb

[E.09a] W. P. E. Book review: Principles of Breeding, by E. Davenport. Biometrika, 6(4):444, March 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331466.

E:1909:BRBa

[E.09b] W. P. E. Book review: The Force of Adverse Selection among Entrants at the Extremes of Life, by Eugene L. Fisk. Biometrika, 6(4):444, March 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331465.

Edgeworth:1907:SOW

[Edg07] F. Y. Edgeworth. Statistical observations on wasps and bees. Biometrika, 5(4):365–386, June 1907. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331685.

Edwards:1908:VDG

[Edw08] Charles Lincoln Edwards. Variation, development and growth in Holothuria Floridana Pourtales and in Holothuria Atra Jager. Biometrika, 6(2/3):236–301, September 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331473.

Elderton:1902:IFD

[Eld02a] W. Palin Elderton. Interpolation by finite differences. Biometrika, 2(1):105–108, November 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331503.

Elderton:1902:TTG

[Eld02b] W. Palin Elderton. Tables for testing the goodness of fit of theory to observation. Biometrika, 1(2):155–163, January 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331485.

Elderton:1903:AGA

[Eld03a] W. Palin Elderton. Addendum to “graduation and analysis of a sickness table. Biometrika, 2(4):503–504, November 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331512.

Elderton:1903:GAS

[Eld03b] W. Palin Elderton. Graduation and analysis of a sickness table. Biometrika, 2(3):260–272, June 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331602.

Elderton:1903:TPN

[Eld03c] W. Palin Elderton. Tables of powers of natural numbers and of the sums of powers of the natural numbers from 1–100. Biometrika, 2(4):474–480, November 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331509.

Elderton:1905:NSP

[Eld05] W. Palin Elderton. Notes on statistical processes. Biometrika, 4(3):374–384, November 1905. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331620.

Elderton:1908:SNI

[Eld08] W. Palin Elderton. Some notes on interpolation in n-dimension space. Biometrika, 6(1):94–103, March 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331558.

Elderton:1909:ADO

[Eld09] Ethel M. Elderton. On the association of drawing with other capacities in school children. Biometrika, 7(1/2):222–226, July/October 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2345372.

Fawcett:1902:SSV

[FL02] Cicely D. Fawcett and Alice Lee. A second study of the variation and correlation of the human skull, with special reference to the Naqada crania. Biometrika, 1(4):408–467, August 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331629.

Fry:1902:NVL

[Fry02] Agnes Fry. Note on variation in leaves of mulberry trees. Biometrika, 1(2):258–259, January 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331491.

Gain:1904:EBV

[Gai04] Edmond Gain. Étude biométrique sur les variations de la fleur et sur l’hétérostylie de Pulmonaria Officinalis L.. (French) [Biometric study on variations of the flower and the heterostyly of Pulmonaria officinalis L.]. Biometrika, 3(4):398–458, November 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331730.

Galton:1901:B

[Gal01] Francis Galton. Biometry. Biometrika, 1(1):7–10, October 1901. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331669.

Galton:1907:GDI

[Gal07] Francis Galton. Grades and deviates: (including a table of normal deviates corresponding to each millesimal grade in the length of an array, and a figure). Biometrika, 5(4):400–406, June 1907. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331687.

Galloway:1909:CBP

[Gal09] A. Rudolf Galloway. Canary breeding. A partial analysis of records from 1891–1909. Biometrika, 7(1/2):1–42, July/October 1909. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2345361.

Griffiths:1904:MOH

[GD04] G. B. Griffiths and H. B. Donkin. Measurements of one hundred and thirty criminals. Biometrika, 3(1):60–62, January 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331522.

Gibson:1906:TFC

[Gib06] Winifred Gibson. Tables for facilitating the computation of probable errors. Biometrika, 4(4):385–393, March 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331636.

Gladstone:1905:SRB

[Gla05] Reginald J. Gladstone. A study of the relations of the brain to the size of the head. Biometrika, 4(1/2):105–123, June 1905. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331533.

Goring:1909:BRB

[Gor09] Charles Goring. Book review: Les femmes homicides. Biometrika, 7 (1/2):231–233, July/October 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2345374.

Greenwood:1904:FSW

[Gre04] M. Greenwood. A first study of the weight, variability, and correlation of the human viscera, with special reference to the healthy and diseased heart. Biometrika, 3(1):63–83, January 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331523.

Greenwood:1909:BSP

[GW09] M. Greenwood, Junr. and J. D. C. White. A biometric study of phagocytosis with special reference to the “opsonic index”: First memoir. On the frequency distributions of phagocytic counts. Biometrika, 6(4):376–401, March 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331459.

Harris:1909:CBV

[Har09a] J. Arthur Harris. The correlation between a variable and the deviation of a dependent variable from its probable value. Biometrika, 6(4):438– 443, March 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331464.

Harris:1909:NVA

[Har09b] J. Arthur Harris. Note on variation in adoxa. Biometrika, 7 (1/2):218–222, July/October 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2345371.

Harris:1909:SMC

[Har09c] J. Arthur Harris. A short method of calculating the coefficient of correlation in the case of integral variates. Biometrika, 7(1/2):214–218, July/October 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2345370.

Hensgen:1902:BUS

[Hen02] C. Hensgen. Biometrische Untersuchungen über die Spielarten von Helix Nemoralis. (German) [Biometric studies on the varieties of Helix nemoralis]. Biometrika, 1(4):468–492, August 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331630.

Heron:1906:ISR

[Her06] David Heron. On the inheritance of the sex-ratio. Biometrika, 5(1/2):79–85, October 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331648.

Heron:1908:ISR

[Her08] David Heron. Inheritance of the sex-ratio in the thoroughbred racehorse. Biometrika, 6(1):120–122, March 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331567.

Heyer:1909:LBG

[Hey09] A. Heyer. Über die Langenvariation der Coniferennadeln. (German) [On the length variation of Coniferennadeln]. Biometrika, 6(4):354– 365, March 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331457.

Harvey:1909:OIM

[HM09] W. F. Harvey and Anderson McKendrick. The opsonic index — a medico-statistical enquiry. Biometrika, 7(1/2):64–95, July/October 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2345364.

Jenkinson:1906:RBS

[Jen06] J. W. Jenkinson. On the relation between the symmetry of the egg and the symmetry of the embryo in the frog (rana temporaria). Biometrika, 5(1/2):147–167, October 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331651.

Jenkinson:1909:RBS

[Jen09] J. W. Jenkinson. On the relation between the symmetry of the egg, the symmetry of segmentation, and the symmetry of the embryo in the frog. Second communication. Biometrika, 7(1/2):148–209, July/ October 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2345368.

Jacob:1903:CNP

[JLP03] S. Jacob, A. Lee, and Karl Pearson. Craniological notes: Preliminary note on interracial characters and their correlation in man. Biometrika, 2(3):347–356, June 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331612.

Kribs:1908:NRV

[Kri08] H. G. Kribs. Note on the relative variability of the sexes in Carabus Auratus, L. Biometrika, 6(1):103–105, March 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331559.

Latter:1902:EBC

[Lat02] Oswald H. Latter. The egg of Cuculus Canorus. An enquiry into the dimensions of the Cuckoo’s egg and the relation of the variations to the size of the eggs of the foster-parent, with notes on coloration, &c. Biometrika, 1(2):164–176, January 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331486.

Latter:1905:EBC

[Lat05] Oswald H. Latter. The egg of Cuculus Canorus: An attempt to ascertain from the dimensions of the Cuckoo’s egg if the species is tending to break up into sub-species, each exhibiting a preference for some one foster-parent: Second memoir. Biometrika, 4(3):363–373, November 1905. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331619.

Lewis:1908:SHS

[LE08] Thomas Lewis and Dennis Embleton. Split-hand and split-foot deformities, their types, origin, and transmission. Biometrika, 6(1):26–58, March 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331555.

Lee:1902:DLV

[Lee02] Alice Lee. Dr Ludwig on variation and correlation in plants. Biometrika, 1(3):316–319, April 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331544.

Lewis:1908:AMS

[Lew08a] Thomas Lewis. Addendum to memoir: “Split-Hand and Split-Foot Deformities”. Biometrika, 6(1):117–118, March 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331565.

Lewis:1908:NIB

[Lew08b] Thomas Lewis. Note on inheritance of Brachydactyly. Biometrika, 6(2/3):327, September 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331476.

Lewenz:1904:MIC

[LP04] M. A. Lewenz and Karl Pearson. On the measurement of internal capacity from cranial circumferences. Biometrika, 3(4):366–397, November 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331729.

Ludwig:1901:VPM

[Lud01] Prof. Dr. F. Ludwig. Variationsstatistische Probleme und Materialien. (German) [Variational-statistical problems and materials].

Biometrika, 1(1):11–29, October 1901. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331670.

Lutz:1903:NIC

[Lut03] Frank E. Lutz. Note on the influence of change in sex on the intensity of heredity. Biometrika, 2(2):237–240, February 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331684.

Lewenz:1902:DPE

[LW02] M. A. Lewenz and M. A. Whiteley. Data for the problem of evolution in man. A second study of the variability and correlation of the hand. Biometrika, 1(3):345–360, April 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331547.

Macdonell:1902:CAI

[Mac02a] W. R. Macdonell. On criminal anthropometry and the identification of criminals. Biometrika, 1(2):177–227, January 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331487.

Macdonell:1902:IPV

[Mac02b] W. R. Macdonell. On the influence of previous vaccination in cases of smallpox. Biometrika, 1(3):375–383, April 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331550.

MacDonell:1903:FSS

[Mac03] W. R. MacDonell. A further study of statistics relating to vaccination and smallpox. Biometrika, 2(2):135–144, February 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331678.

MacDonell:1904:SVC

[Mac04] W. R. MacDonell. A study of the variation and correlation of the human skull, with special reference to English crania. Biometrika, 3(2/3):191–244, March/July 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331480.

Macdonell:1906:SSE

[Mac06] W. R. Macdonell. A second study of the English skull, with special reference to Moorfields crania. Biometrika, 5(1/2):86–104, October 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331649.

Maynard:1909:SSA

[May09] G. D. Maynard. Statistical study of anti-typhoid inoculation. Biometrika, 6(4):366–375, March 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331458.

McIntosh:1903:VBN

[McI03] D. C. McIntosh. Variation in “Ophiocoma Nigra” (O. F. Muller). Biometrika, 2(4):463–473, November 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331508.

Myers:1903:CNH

[Mye03] Charles S. Myers. Craniological notes: Homogeneity and heterogeneity in crania. Biometrika, 2(4):504–505, November 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331513. See remarks [P.03a].

P:1901:SBM

[P.01] K. P. Statoblasts of Pectinatella Magnifica. Biometrika, 1(1):128, October 1901. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331676.

P:1903:CNRa

[P.03a] K. P. Craniological notes: Remarks on Dr. C. S. Myers’ note. Biometrika, 2(4):506–508, November 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331514. See [Mye03].

P:1903:CNRb

[P.03b] K. P. Craniological notes: Remarks on Professor Aurel von Torok’s note. Biometrika, 2(4):509–512, November 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331516.

P:1904:MVL

[P.04a] K. P. A Mendelian’s view of the law of ancestral inheritance. Biometrika, 3(1):109–112, January 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331528.

P:1904:CBH

[P.04b] K. P. On the correlation between hair colour and eye colour in man. Biometrika, 3(4):459–462, November 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331731.

P:1907:BRBa

[P.07a] K. P. Book review: On the Theory of Inheritance of Quantitatively Compound Characters on the Basis of Mendel’s Laws, by G. Udny Yule. Biometrika, 5(4):481–482, June 1907. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331701.

P:1907:BRBb

[P.07b] K. P. Book review: Über Dolichocephalie und Brachycephalie. Zur Kritik der Index-Angaben. by W. Johannsen. Biometrika, 5(4):482, June 1907. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331702.

P:1908:NIM

[P.08] K. P. Note on inheritance in man. Biometrika, 6(2/3):327–328, September 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331477.

P:1909:BRB

[P.09] R. P. Book review: Mental and Moral Heredity in Royalty. A Statistical Study in History and Psychology, by Frederick Adams Woods. Biometrika, 6(4):445, March 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331468.

Pearl:1903:VCB

[PD03] Raymond Pearl and Frances J. Dunbar. Variation and correlation in Arcella. Biometrika, 2(3):321–337, June 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331608.

Pearson:1902:CEL

[Pea02a] K. Pearson. On the change in expectation of life in man during a period of circa 2000 years. Biometrika, 1(2):261–264, January 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331493.

Pearson:1902:MVO

[Pea02b] K. Pearson. On the modal value of an organ or character. Biometrika, 1(2):260–261, January 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331492.

Pearson:1902:VES

[Pea02c] K. Pearson. Variation of the egg of the sparrow (Passer Domesticus). Biometrika, 1(2):256–257, January 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331490.

Pearson:1902:FCB

[Pea02d] Karl Pearson. On the fundamental conceptions of biology. Biometrika, 1(3):320–344, April 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331546.

Pearson:1902:SFCb

[Pea02e] Karl Pearson. On the systematic fitting of curves to observations and measurements. Biometrika, 1(3):265–303, April 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331540.

Pearson:1902:SFCa

[Pea02f] Karl Pearson. On the systematic fitting of curves to observations and measurments: Part II. Biometrika, 2(1):1–23, November 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331498.

Pearson:1903:CNH

[Pea03a] Karl Pearson. Craniological notes: Homogeneity and heterogeneity in collections of crania. Biometrika, 2(3):345–347, June 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331611.

Pearson:1903:CNP

[Pea03b] Karl Pearson. Craniological notes: Professor Aurel von Torok’s attack on the arithmetical mean. Biometrika, 2(3):339–345, June 1903. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331610.

Pearson:1904:NMP

[Pea04a] Karl Pearson. Note on Mr Punnett’s section on the inheritance of meristic characters. Biometrika, 3(4):363–365, November 1904. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331728.

Pearson:1904:LIM

[Pea04b] Karl Pearson. On the laws of inheritance in man: II. On the inheritance of the mental and moral characters in man, and its comparison with the inheritance of the physical characters. Biometrika, 3(2/3):131–190, March/July 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331479.

Pearl:1905:BSM

[Pea05a] Raymond Pearl. Biometrical studies on man: I. Variation and correlation in brain-weight. Biometrika, 4(1/2):13–104, June 1905. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331532.

Pearson:1905:FSV

[Pea05b] Karl Pearson. “Das Fehlergesetz und Seine Verallgemeinerungen Durch Fechner und Pearson.” A rejoinder. Biometrika, 4(1/2):169– 212, June 1905. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331536.

Pearl:1906:CPE

[Pea06a] Raymond Pearl. The calculation of the probable errors of certain constants of the normal curve. Biometrika, 5(1/2):190, October 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331659.

Pearl:1906:MDL

[Pea06b] Raymond Pearl. On the mean duration of life of individuals dying within a year after birth. Biometrika, 4(4):510–516, March 1906. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331642.

Pearl:1906:VBU

[Pea06c] Raymond Pearl. Variation in Chilomonas under favourable and unfavorable conditions. Biometrika, 5(1/2):53–72, October 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331646.

Pearson:1906:NSN

[Pea06d] Karl Pearson. Note on the significant or non-significant character of a sub-sample drawn from a sample. Biometrika, 5(1/2):181–183, October 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331656.

Pearson:1906:CCH

[Pea06e] Karl Pearson. On a coefficient of class heterogeneity or divergence. Biometrika, 5(1/2):198–203, October 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331661.

Pearson:1906:CPC

[Pea06f] Karl Pearson. On certain points connected with scale order in the case of the correlation of two characters which for some arrangement give a linear regression line. Biometrika, 5(1/2):176–178, October 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331654.

Pearson:1906:CWM

[Pea06g] Karl Pearson. On the curves which are most suitable for describing the frequency of random samples of a population. Biometrika, 5(1/2):172– 175, October 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331653.

Pearson:1906:DSA

[Pea06h] Karl Pearson. On the distribution of severity of attack in cases of smallpox. Biometrika, 4(4):505–510, March 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331641. See [Pea07b, Tur07].

Pearson:1906:RIS

[Pea06i] Karl Pearson. On the relationship of intelligence to size and shape of head, and to other physical and mental characters. Biometrika, 5(1/2):105–146, October 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331650.

Pearson:1906:RPK

[Pea06j] Karl Pearson. A rejoinder to Professor Kapteyn. Biometrika, 5(1/2): 168–169+171+171, October 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331652.

Pearl:1907:BSC

[Pea07a] Raymond Pearl. A biometrical study of conjugation in paramecium. Biometrika, 5(3):213–297, February 1907. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331631.

Pearson:1907:FRD

[Pea07b] Karl Pearson. Further remarks on the distribution of severity in cases of smallpox. Biometrika, 5(3):362–364, February 1907. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331635. See [Pea06h, Tur07].

Pearson:1907:RCC

[Pea07c] Karl Pearson. Reply to certain criticisms of Mr G. U. Yule. Biometrika, 5(4):470–476, June 1907. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331693.

Pearl:1908:CPC

[Pea08a] Raymond Pearl. On certain points concerning the probable error of the standard deviation. Biometrika, 6(1):112–117, March 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331564.

Pearson:1908:HSR

[Pea08b] Karl Pearson. On heredity in sex. Remarks on Mr Cobb’s note. Biometrika, 6(1):109–111, March 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331562.

Pearson:1908:IDK

[Pea08c] Karl Pearson. On inheritance of the deformity known as split-foot or lobster-claw. Biometrika, 6(1):69–93, March 1908. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331557.

Pearson:1908:IDS

[Pea08d] Karl Pearson. On the influence of double selection on the variation and correlation of two characters. Biometrika, 6(1):111–112, March 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331563.

Pearl:1909:FCV

[Pea09a] Raymond Pearl. The frequency constants of a variable z = f(x_{1}, x_{2}). Biometrika, 6(4):437–438, March 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331463.

Pearson:1909:NSC

[Pea09b] Karl Pearson. Note on the skin-colour of the crosses between Negro and White. Biometrika, 6(4):348–353, March 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331456.

Pearson:1909:NMD

[Pea09c] Karl Pearson. On a new method of determining correlation between a measured character A, and a character B, of which only the percentage of cases wherein B exceeds (or falls short of) a given intensity is recorded for each grade of A. Biometrika, 7(1/2):96–105, July/ October 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2345365.

Pearson:1909:BSR

[Pea09d] Karl Pearson, D.Sc. A biometric study of the red blood corpuscles of the common tadpole (Rana Temporaria), from the measurements of Ernest Warren. Biometrika, 6(4):402–419, March 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331460.

Perrin:1904:SDE

[Per04a] Emily Perrin. On some dangers of extrapolation. Biometrika, 3(1):99– 103, January 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331525.

Perrin:1904:CBO

[Per04b] Emily Perrin. On the contingency between occupation in the case of fathers and sons. Biometrika, 3(4):467–469, November 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331733.

Pearl:1905:VCE

[PF05] Raymond Pearl and Wilbur N. Fuller. Variation and correlation in the earthworm. Biometrika, 4(1/2):213–229, June 1905. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331537.

Pearson:1903:LIM

[PL03] Karl Pearson and Alice Lee. On the laws of inheritance in man: I. Inheritance of physical characters. Biometrika, 2(4):357–462, November 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331507.

Pearson:1908:GPE

[PL08] Karl Pearson and Alice Lee. On the generalised probable error in multiple normal correlation. Biometrika, 6(1):59–68, March 1908. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331556.

Powys:1901:DPE

[Pow01] A. O. Powys. Data for the problem of evolution in man. Anthropometric data from Australia. Biometrika, 1(1):30–49, October 1901. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331671.

Powys:1905:DPE

[Pow05] A. O. Powys. Data for the problem of evolution in man. On fertility, duration of life and reproductive selection. Biometrika, 4(3):233–285, November 1905. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331614.

Pearl:1909:DVC

[PP09] Raymond Pearl and Maud Dewitt Pearl. Data on variation in the comb of the domestic fowl. Biometrika, 6(4):420–432, March 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331461.

Pearson:1904:DHL

[PR04] Karl Pearson and Marion Radford. On differentiation and homotyposis in the leaves of Fagus Sylvatica. Biometrika, 3(1):104–107, January 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331526.

Punnett:1904:MSS

[Pun04] R. C. Punnett. Merism and sex in “spinax niger”. Biometrika, 3(4):313–362, November 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331727.

Pearson:1902:VRF

[PY02] K. Pearson and G. U. Yule. Variation in ray-flowers of Chrysanthemum Leucanthemum, 1133 heads gathered at Keswick, during July, 1895, by K. Pearson and G. U. Yule. Biometrika, 1(3):319, April 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331545.

Pearson:1903:LAH

[PYBL03] Karl Pearson, G. U. Yule, Norman Blanchard, and Alice Lee. The law of ancestral heredity. Biometrika, 2(2):211–236, February 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331683.

Rhind:1909:TFC

[Rhi09] A. Rhind. Tables to facilitate the computation of the probable errors of the Chief constants of skew frequency distributions. Biometrika, 7 (1/2):127–147, July/October 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2345367.

Rietz:1909:IPB

[Rie09] H. L. Rietz. On inheritance in the production of butter fat. Biometrika, 7(1/2):106–126, July/October 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2345366.

Robbins:1908:VFH

[Rob08] W. W. Robbins. Variation in flower-heads of Gaillardia Aristata. Biometrika, 6(1):106–108, March 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331560.

Rommel:1906:IFL

[RP06] G. M. Rommel and E. F. Phillips. Inheritance in the female line of size of litter in Poland China sows. Biometrika, 5(1/2):203–205, October 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331662.

S:1902:MSP

[S02] Francis Galton F. R. S. The most suitable proportion between the value of first and second prizes. Biometrika, 1(4):385–399, August 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331627.

Schuster:1903:VBP

[Sch03] E. H. J. Schuster. Variation in “Eupagurus Prideauxi” (Heller). Biometrika, 2(2):191–210, February 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331682.

Schuster:1905:LBR

[Sch05a] E. H. J. Schuster. The long Barrow and round Barrow skulls in the collection of the department of comparative anatomy, the museum, Oxford. Biometrika, 4(3):351–362, November 1905. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331618.

Schuster:1905:RCG

[Sch05b] E. H. J. Schuster. Results of crossing grey (house) mice with albinos. Biometrika, 4(1/2):1–12, June 1905. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331531.

Schuster:1906:HDD

[Sch06a] Edgar Schuster. Hereditary deafness: A discussion of the data collected by Dr E. A. Fay in America. Biometrika, 4(4):465–482, March 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331639.

Schuster:1906:PZG

[Sch06b] Edgar Schuster. Professor Ziegler and Galton’s Law of Ancestral Inheritance. Biometrika, 5(1/2):184, October 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331657.

Schtscherbakow:1908:FVT

[Sch08] Th. S. Schtscherbakow. Zur Frage Vom Viergliederigen Tarsus Der Blattidae Und Der Regeneration Der Füsse Derselben. (German) [On the question of the quadrapartite tarsus of Blattidae regeneration and the feet of the same]. Biometrika, 6(2/3):311–326, September 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331475.

Schuster:1907:IPC

[SE07] Edgar Schuster and E. M. Elderton. The inheritance of psychical characters. Biometrika, 5(4):460–469, June 1907. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331692.

Sheppard:1903:NTP

[She03] W. F. Sheppard. New tables of the probability integral. Biometrika, 2(2):174–190, February 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331681.

Sheppard:1907:CMF

[She07] W. F. Sheppard. The calculation of moments of a frequencydistribution. Biometrika, 5(4):450–459, June 1907. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331691.

Simpson:1902:RBF

[Sim02] J. Y. Simpson. The relation of binary fission to variation. Biometrika, 1(4):400–407, August 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331628.

Smith:1903:BEB

[Smi03] Geoffrey Smith. Actinosphaerium Eichorni: A biometrical study in the mass relations of nucleus and cytoplasm. Biometrika, 2(3):241– 254, June 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331600.

Sommerville:1906:CFR

[Som06] D. M. Y. Sommerville. On the classification of frequency-ratios. Biometrika, 5(1/2):179–181, October 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331655.

Spencer:1904:GST

[Spe04] John Spencer. Graduation of a sickness table by Makeham’s hypothesis. Biometrika, 3(1):52–57, January 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331520.

Student:1907:ECH

[Stu07] Student. On the error of counting with a haemacytometer. Biometrika, 5(3):351–360, February 1907. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331633.

Student:1908:PEC

[Stu08a] Student. Probable error of a correlation coefficient. Biometrika, 6 (2/3):302–310, September 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331474.

Student:1908:PEM

[Stu08b] Student. The probable error of a mean. Biometrika, 6(1):1–25, March 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331554.

Student:1909:DMS

[Stu09] Student. The distribution of the means of samples which are not drawn at random. Biometrika, 7(1/2):210–214, July/October 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2345369.

Surface:1909:FS

[Sur09] Frank M. Surface. Fecundity of swine. Biometrika, 6(4):433–436, March 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331462.

Thomson:1909:SCS

[TBP09] E. Y. Thomson, Julia Bell, and Karl Pearson. A second cooperative study of Vespa vulgaris. Comparison of queens of a single nest and queens of a general population. Biometrika, 7(1/2):48–63, July/ October 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2345363.

Tocher:1906:AM

[Toc06a] J. F. Tocher. Appendix 1: Measurements. Biometrika, 5(supplement): 5–80, ???? 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331579.

Anthropometric survey of the inmates of the asylums in Scotland.

Tocher:1906:ATC

[Toc06b] J. F. Tocher. Appendix 2: Tables of classified data. Biometrika, 5(supplement):81–96, ???? 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2331580. Anthropometric survey of the inmates of the asylums in Scotland.

Tocher:1907:ACI

[Toc07] J. F. Tocher. The anthropometric characteristics of the inmates of asylums in Scotland. Biometrika, 5(3):298–350, February 1907. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331632.

Tocher:1908:PSS

[Toc08] J. F. Tocher. Pigmentation survey of school children in Scotland. Biometrika, 6(2/3):129–235, September 1908. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331472.

Tocher:1909:PSS

[Toc09] J. F. Tocher. Pigmentation survey of school children in Scotland. Biometrika, 6(4):1–67, March 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331470.

Tower:1902:VRF

[Tow02] W. L. Tower. Variation in the ray-flowers of Chrysanthemum Leucanthemum L. at Yellow Springs, Greene Co., O., with remarks upon the determination of modes. Biometrika, 1(3):309–315, April 1902. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331543.

Tschepourkowsky:1905:CSI

[Tsc05a] E. Tschepourkowsky. Contributions to the study of interracial correlation. Biometrika, 4(3):286–312, November 1905. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331615.

Tschepourkowsky:1905:QSR

[Tsc05b] E. Tschepourkowsky. A quantitative study of the resemblance between man and woman. Biometrika, 4(1/2):161–168, June 1905. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331535.

Turner:1906:CBV

[Tur06] F. M. Turner. On the correlation between vaccination and smallpox in the London epidemic, 1901–2. Biometrika, 4(4):483–504, March 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331640.

Turner:1907:DSA

[Tur07] F. M. Turner. On the distribution of severity of attack in cases of smallpox. Biometrika, 5(3):361–362, February 1907. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331634.

vonTorok:1903:CNN

[vT03] Aurel von Torok. Craniological notes: Note on cranial types. Biometrika, 2(4):508–509, November 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331515.

W:1902:SCC

[W.02] W. F. R. W. Seasonal change in the characters of Aster Prenanthoides, Muhl. Biometrika, 2(1):113–114, November 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331505.

W:1903:IPV

[W.03] W. F. R. W. Inheritance in phaseolus vulgaris. Biometrika, 2 (4):499–503, November 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331511.

W:1904:ASM

[W.04] W. F. R. W. Albinism in Sicily and Mendel’s laws. Biometrika, 3(1):107–109, January 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2331527. See corrections [Bat04, Bat05].

Warren:1902:VIP

[War02] Ernest Warren. Variation and inheritance in the parthenogenetic generations of the aphis “Hyalopterus Trihodus” (Walker). Biometrika, 1(2):129–154, January 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331484.

Warren:1903:PAA

[War03] Ernest Warren. A preliminary attempt to ascertain the relationship between the size of cell and the size of body in Daphnia Magna Straus. Biometrika, 2(3):255–259, June 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331601.

Warren:1909:SSO

[War09] Ernest Warren. Some statistical observations on termites, mainly based on the work of the late Mr G. D. Haviland. Biometrika, 6(4):329–347, March 1909. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331455.

Weldon:1901:COC

[Wel01a] W. F. R. Weldon. Change in organic correlation of ficaria ranunculoides during the flowering season. Biometrika, 1(1):125–128, October 1901. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2331675.

Weldon:1901:FSN

[Wel01b] W. F. R. Weldon. A first study of natural selection in Clausilia Laminata (Montagu). Biometrika, 1(1):109–124, October 1901. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http:

//www.jstor.org/stable/2331674.

Weldon:1902:MLA

[Wel02a] W. F. R. Weldon. Mendel’s laws of alternative inheritance in peas. Biometrika, 1(2):228–254, January 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331488.

Weldon:1902:AMC

[Wel02b] W. F. R. Weldon. On the ambiguity of Mendel’s categories. Biometrika, 2(1):44–55, November 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331500.

Weldon:1902:PVO

[Wel02c] W. F. R. Weldon. Professor de Vries on the origin of species. Biometrika, 1(3):365–374, April 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331549.

Weldon:1903:MBR

[Wel03] W. F. R. Weldon. Mr Bateson’s revisions of Mendel’s theory of heredity. Biometrika, 2(3):286–298, June 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331605.

Weldon:1904:NRB

[Wel04] W. F. R. Weldon. Note on a race of Clausilia Itala (Von Martens). Biometrika, 3(2/3):299–307, March/July 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331482.

Weldon:1907:HMR

[Wel07] W. F. R. Weldon. On heredity in mice from the records of the late W. F. R. Weldon. On the inheritance of the sex-ratio and of the size of litter. Biometrika, 5(4):436–449, June 1907. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331690.

Whitehead:1902:VM

[Whi02] Henry Whitehead. Variation in the moscatel. Biometrika, 2 (1):108–113, November 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331504.

Wright:1907:CSQ

[WLP07] Alexandra Wright, Alice Lee, and K. Pearson. A cooperative study of queens, drones and workers in “Vespa Vulgaris”. Biometrika, 5(4): 407–422, June 1907. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2331688.

Woods:1903:MLS

[Woo03] Frederick Adams Woods. Mendel’s laws and some records in rabbit breeding. Biometrika, 2(3):299–306, June 1903. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331606.

Woods:1906:NIS

[Woo06] Frederick Adams Woods. The non-inheritance of sex in man. Biometrika, 5(1/2):73–78, October 1906. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2331647.

Yule:1902:LDR

[Yul02a] G. U. Yule. Local death rates. Biometrika, 1(3):384, April 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331551.

Yule:1902:VNS

[Yul02b] G. Udny Yule. Variation of the number of sepals in Anemone Nemorosa. Biometrika, 1(3):307–309, April 1902. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331542.

Yule:1903:NTA

[Yul03] G. Udny Yule. Notes on the theory of association of attributes in statistics. Biometrika, 2(2):121–134, February 1903. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331677.

Yule:1904:CMD

[Yul04] G. Udny Yule. On a convenient means of drawing curves to various scales. Biometrika, 3(4):469–471, November 1904. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331734.